### v1

# A degree-biased cutting process for random recursive trees

Abstract
We study a degree-biased cutting process on random recursive trees, where vertices are deleted with probability proportional to their degree. We verify the splitting property and explicitly obtain the distribution of the number of vertices deleted by each cut. This allows us to obtain a recursive formula for $K_{n}$, the number of cuts needed to destroy a random recursive tree of size $n$. Furthermore, we show that $K_{n}$ is stochastically dominated by $J_{n}$, the number of jumps made by a certain random walk with a barrier. We obtain a convergence in distribution of $J_{n}$ to a Cauchy random variable. We also explore the relationship of this cutting procedure to a coalescing process and show that the coalescing rate cannot correspond to a $\Lambda$-coalescent since it fails consistency conditions.
1 Introduction
An increasing tree on $[n]$ is a rooted labeled tree, where labels along a path away from the root are in increasing order. We write $I_{n}$ for the set of increasing trees of size $n$; namely, with $n$ vertices. A random recursive tree on $[n]$, denoted $T_{n}$ is a tree chosen uniformly at random from $I_{n}$. A classical construction of $T_{n}$, for $n\geq 1$, is to start with a vertex $1$ as the root and for $2\leq k\leq n$, randomly choose a vertex from $\{1,\dots,k-1\}$ and attach vertex $k$ to the chosen vertex.
The idea of cutting down random networks was introduced in [17] for random rooted labeled trees and soon afterward studied for the case of random recursive trees [18]. In the original procedure, the objective is to disconnect a given vertex (the root) by sequentially removing a uniformly chosen edge and discarding the subtree that does not contain the root. The procedure ends when there is only one vertex left; we say that the root has been isolated or that the tree has been detroyed.
Let $X_{n}$ be the number of cuts necessary to isolate the root of a random recursive tree on $n$ vertices. In [18], Meir and Moon observe that the distribution of the remaining tree after one cut is, conditionally given its size, a random recursive tree. This characteristic of the model was later known as the splitting property. In [19], Panholzer obtains a law of large numbers for $X_{n}$ and, using different methods (singularity analysis of the generating functions and a probabilistic argument reviewed in Section 2), both Drmota et al, and Iksanov and Möhle deduce that the fluctuations of $X_{n}$, after an appropriate normalization, converge in distribution to a Cauchy random variable [9, 15]; see Section 2.1 for detailed expressions of these results.
Several other classes of random trees have been studied; see [2] for a thorough survey on cutting procedures and other related processes on random recursive trees up to 2014. Starting in 2019, Cai et al. propose a cutting procedure that provides a notion for network resilience. In the $k$-cut model, again, the edges are selected uniformly at random but only removed after they have received $k$ cuts [6, 7, 21].
The edge-removal cutting procedure can be modeled by placing exponential clocks on each of the edges of the tree deleting them according to the order induced by the clocks. To count the number of cuts necessary to isolate the root, it suffices to consider clocks/edges that belong to the root component at the moment of ringing. This perspective gives rise to the so-called cut-tree $\mathrm{Cut}(T_{n})$ and an associated coalescent processes; see Section 4.
We note that removing a uniformly chosen edge is equivalent to removing a vertex uniformly chosen from the current vertex set except the root. Thus, the vertex-removal process is asymptotically equivalent to edge-removal procedures. The choice of model is determined by the splitting properties of the random trees under study.
Dieuleveut studied a degree-biased vertex-removal process for Galton-Watson trees with offspring distribution belonging to the domain of attraction of a stable law of index $\alpha\in(1,2)$ [8]; namely, a vertex is selected with probability proportional to its degree, then all its descendants (and their subtrees) are removed but the selected vertex is kept. The results on the cut-trees for Galton-Watson trees with finite variance are extended; namely, the corresponding rescaled cut-trees converge in distribution to the stable tree of index $\alpha$, in the sense induced by the Gromov–Prokhorov topology. From the rescaling factor in this setting, it can be inferred that the number of cuts required to isolate the root has first order $n^{1-1/\alpha}$.
The goal of this paper is to devise cutting procedures for random recursive trees that speed up the destruction of the tree, meaning that fewer cuts are needed to isolate the root. Our main results are concerned with $K_{n}$, the number of cuts required to destroy a random recursive tree $T_{n}$ on $n$ vertices in the following process.
The degree-biased cutting procedure, an iterative process of degree-biased vertex-removals until the root is isolated or the tree is destroyed instantly. In either case, we say that the tree has been destroyed (see Definitions 1.1 and 1.4). This vertex-removal process is similar to the one studied by Dieuleveut in the sense that the probability a vertex is selected in the process is proportional to its degree. The main difference is that our model deletes the selected vertex, instead of removing only their offspring. The choice of removing the selected vertex allows us to maintain the splitting property, (see Proposition 1.2), which is key to the probabilistic technique on establishing the limiting distribution of $X_{n}$. In fact, such approach was later studied in [16] for jump distributions satisfying more general conditions; in particular, the number of coalitions in a beta-coalescent until coagulation.
Although the splitting property is maintained, the techniques used in [15, 16] are not straightforward to apply to the degree-biased deletion process, see Remark 3.3. In the case of degree-biased cuts to $T_{n}$, it is only possible to approximate the number of vertices removed with a jump distribution $\zeta$, from which we obtain a stochastic upper bound for $K_{n}$ whose asymptotic growth is known explicitly. More specifically, Theorem 1.5 together with Theorem 1.6 show that the growth order of $K_{n}$ is faster than that of $X_{n}$ at least by a factor of $\ln n$.
In the same spirit, the targeted cutting process for random recursive trees consists of a deterministic procedure (conditionally given the degree sequence of the tree $T_{n}$) that dramatically reduces the necessary number of cuts required to isolate the root [11]. More precisely, the process starts by listing the vertices in decreasing order of their degrees and proceeds with a vertex deletion in the listed order. See Section 1.2 for a comparison of the results.
1.1 Definition of the degree-biased cutting process
In what follows it is relevant to make a difference between random recursive trees $T_{n}$ and their deterministic feasible values; thus, deterministic trees will be denoted with lowercase letters; e.g., $t\in I_{n}$. We consider edges in increasing trees as being directed towards the root. In particular, for an increasing tree $t$ and vertex $v\in t$, we define the degree of $v$, denoted $\deg_{t}(v)$, as the number of edges directed towards $v$. If $v_{1}v_{2}$ is an edge of an increasing tree, we will implicitly assume that $v_{1}<v_{2}$. Whenever $v_{1}>1$, there is a unique outgoing edge from $v_{1}$ to a vertex (its parent) which we denote $v_{0}$.
Let $t^{\prime}$ be a tree with vertex set $\{v_{1},\dots,v_{m}\}$, where $v_{1}<\dots<v_{m}$. We denote $\Phi_{m}$ to the function that relabels each $v_{i}\in t^{\prime}$ with label $i$. In particular, if $t\in I_{n}$ and $t^{\prime}\subset t$ has size $m$ then $\Phi(t^{\prime})\in I_{m}$. In what follows, we abuse notation and write $\Phi(\cdot)=\Phi_{m}(\cdot)$ whenever the size of the tree is clear from the context. Finally, for a vertex $v\in t$, let $t^{(v)}$ denote the subtree rooted at $v$; i.e. $t^{(v)}\subset t$ contains $v$ and all vertices $w$ such that there is a directed path from $w$ to $v$. The following definition of degree-biased cutting has the property that, at each step, the probability of deleting vertex $v$ is proportional to its degree.
Definition 1.1 (Degree-biased cut).
Let $t$ be an increasing tree. Choose an edge $v_{1}v_{2}$ uniformly at random and delete vertex $v_{1}$ along with $t^{(v_{1})}$. More precisely, if $v_{1}>1$, then remove the edge $v_{0}v_{1}$ and keep only the subtree containing the root of $t$; that is, keep $t^{root}:=t\setminus t^{(v_{1})}$. Otherwise, $v_{1}$ is the root of $t$, we let $t^{root}:=\emptyset$ and say that the cut destroys the tree instantly.
A key property for our results is that the splitting property holds for the degree-biased cutting of random recursive trees.
Proposition 1.2 (Splitting property).
Let $n\geq 2$ and let $T_{n}$ be a random recursive tree $T_{n}$. Let $T_{n}^{root}$ be the tree obtained after a degree-biased cut. For $\ell\in[n-2]$, conditionally given that $|T_{n}^{root}|=\ell$, $\Phi(T_{n}^{root})$ has the distribution of a random recursive tree of size $\ell$.
Note that $|T_{n}^{(root)}|\leq n-2$ since the cutting process removes at least both ends of the selected edge. Consequently, the size of each cut satisfies $2\leq|T_{n}^{(v_{1})}|\leq n$. Moreover, we obtain its distribution by exploiting the fact that ${\mathbf{E}}\left[\deg_{T_{n}}(1)\right]=H_{n-1}$, where $H_{n}=\sum_{i=1}^{n}i^{-1}$ is the $n$-th harmonic number.
Proposition 1.3 (Size of a cut).
Let $n\geq 2$ and let $T_{n}$ be a random recursive tree $T_{n}$. Let $T_{n}^{(v_{1})}$ be the tree defined by a degree-biased cut of $T_{n}$; where $v_{1}v_{2}$ is a uniformly chosen edge of $T_{n}$. The distribution of $|T_{n}^{(v_{1})}|$ is given by
${\mathbf{P}}\left(|T_{n}^{(v_{1})}|=k\right)=\frac{nH_{k-1}}{(n-1)(k+1)k}{\mathbf{1}}_{[2\leq k<n]}+\frac{H_{n-1}}{n-1}{\mathbf{1}}_{[k=n]}.$ (1)
We note in passing that the probability that the degree-biased cutting procedure is destroyed in one step has a relatively large mass coming from an instant destruction of the tree; see Remark 3.3.
Definition 1.4 (Degree-biased process).
Let $n\geq 2$ and $T_{n}$ be a random recursive tree. Sequentially perform degree-biased cuts until the resulting tree (or empty set) cannot be cut anymore; at which point we say that $T_{n}$ was destroyed. More precisely, let $(G_{t})_{t\geq 0}$ be a sequence of trees such that $G_{0}=T_{n}$ and, for $t\geq 1$, if $|G_{t-1}|\geq 2$, then $G_{t}=G_{t-1}^{root}$; otherwise, $G_{t}=G_{t-1}$. The number of cuts needed to destroy $T_{n}$ is defined by
$K_{n}=\min\{t\geq 1:|G_{t}|\in\{0,1\}\};$
the halting cases correspond to either instantly destroying the tree or isolating the root.
1.2 Main results
Let $\{\xi_{i}:i\in\mathbb{N}\}$ be a sequence of independent copies of a random variable $\xi$ taking positive integer values and $n\in\mathbb{N}$. The random walk with a barrier at $n$, $\{R^{(n)}_{i}:i\in\mathbb{N}_{0}\}$ is defined recursively as $R^{(n)}_{0}=0$ and, for $i\geq 1$,
$R^{(n)}_{i}:=R^{(n)}_{i-1}+\xi_{i}\mathbf{1}_{\{R_{i-1}+\xi_{i}<n\}}.$
Throughout this section, consider $\xi=\zeta$ with mass probability, for $k\geq 2$,
${\mathbf{P}}\left(\zeta=k\right)=\frac{H_{k-1}}{k(k+1)}.$ (2)
Let $J_{n}$ denote the number of jumps of the process $\{R^{(n+1)}_{i}:i\in\mathbb{N}_{0}\}$ when $\xi=\zeta$; that is, $J_{n}:=\sum_{i\geq 0}\mathbf{1}_{\{R^{(n+1)}_{i-1}+\xi_{i}\leq n\}}$. Using results in [16], which are summarized in Section 2, we obtain the asymptotic behavior of its mean and the limiting distribution of $J_{n}$, under an appropriate normalization.
Theorem 1.5.
As $n$ goes to infinity, ${\mathbf{E}}\left[J_{n}\right]\sim 2n(\ln n)^{-2}$ and, in probability,
$\frac{J_{n}}{{\mathbf{E}}\left[J_{n}\right]}\stackrel{{\scriptstyle p}}{{\longrightarrow}}1,$
and
$\frac{J_{n}}{n}(\ln{n})^{3}-\ln{n}$
converges in distribution to a Cauchy random variable.
Unfortunately, Proposition 1.3 implies that the degree-biased cutting places some extra mass of instant destruction at each step of the process, preventing $J_{n}$ from being a good approximation of $K_{n}$; see Remark 3.3. Nevertheless, Theorem 1.5 gives us insight into the deletion time $K_{n}$ of the degree-biased cutting process on random recursive trees. Indeed, Proposition 1.2 allows us to obtain a recursive formula for $K_{n}$ and to compare $K_{n}$ with $J_{n}$, akin to [15, Lemma 1]. Our main result states that $J_{n}$ may be interpreted as a stochastic overcount of $K_{n}$.
Theorem 1.6.
$J_{n}$ stochastically dominates $K_{n}$.
In conclusion, the degree-biased cutting process results in quicker deletion times compared to the random cutting approach, at least by a factor of $\ln n$ in the first order of $X_{n}$; compare Theorem 1.5 with (5) and (6). In contrast, the degree-biased procedure is significantly slower when compared to deleting vertices deterministically based on their ordered degrees, as in the targeted vertex-cutting process where the number of required cuts has asymptotic growth at most $n^{1-\ln 2}$ [11].
Further research into this topic includes obtaining tighter bounds for the deletion time of both the degree-biased cutting process and the targeted vertex-cutting. In addition, it would be interesting to develop a better understanding of the asymptotic properties of the size of the subtrees removed at each step of the degree-biased cutting procedure, as this is key to further understand the properties of natural, related models such as the cut-tree and coalescent process.
The outline for the paper is the following. In Section 2 we review the results for random walks with a barrier at $n$ [16] and place the known properties of $X_{n}$ into such context.
In Section 3 we establish the splitting property and the distribution of the mass lost at each degree-biased cut of the process; Propositions 1.2 and 1.3. To do so, we compute the joint probability of $T^{root}$ and $T^{(v_{1})}$, the tree removed after the first degree-biased cut, see Lemma 3.1; and use a first-step analysis to recover a recursive formula for $K_{n}$, see Remark 3.2. Section 3.1 contains the proofs of Theorems 1.5 and 1.6.
Finally, in Section 4 we briefly explore both the cut-tree and coalescent associated with the degree-biased cutting procedure.
Notation. We use $|A|$ to denote the cardinality of a set $A$. For $n\in\mathbb{N}$ we write $[n]:=\{1,2,\dots,n\}$. $H_{n}=\sum_{i=1}^{n}i^{-1}$ is the $n$-th harmonic number. We denote natural logarithms by $\ln$. For real functions $f,g$ we write $f(x)\sim g(x)$ when $\lim_{x\to\infty}f(x)/g(x)=1$, $f(x)=o(g(x))$ when $\lim_{x\to\infty}f(x)/g(x)=0$ and $f(x)=O(g(x))$ when $|f(x)/g(x)|\leq C$ for some $C>0$. We use $\stackrel{{\scriptstyle p}}{{\longrightarrow}}$ to denote convergence in probability.
2 Jumps for random walks with a barrier
Let $\{\xi_{i}:i\in\mathbb{N}\}$ be a sequence of independent copies of a random variable $\xi$ taking positive integer values and $n\in\mathbb{N}$. The random walk with a barrier at $n$, $\{R^{(n)}_{i}:i\in\mathbb{N}_{0}\}$ is defined recursively as $R^{(n)}_{0}=0$ and, for $i\geq 1$,
$R^{(n)}_{i}:=R^{(n)}_{i-1}+\xi_{i}\mathbf{1}_{\{R_{i-1}+\xi_{i}<n\}}.$
The process $\{R^{(n)}_{i}:i\in\mathbb{N}_{0}\}$ has non-decreasing paths and satisfies $R_{i}^{(n)}<n$ for all $i\in\mathbb{N}_{0}$. Let $M_{n}=|\{i\in\mathbb{N}|R^{(n)}_{i-1}\neq R^{(n)}_{i}\}|=\sum_{i\geq 0}\mathbf{1}_{\{R^{(n)}_{i-1}+\xi_{i}<n\}}$ denote the number of jumps of the process $\{R^{(n)}_{i}:i\in\mathbb{N}_{0}\}$. Note that $M_{n}\leq n$ since $\xi\geq 1$ with probability one.
The asymptotic behavior of $M_{n}$ has been studied when $\xi$ is distributed as in (4) in [15] and for broader classes of $\xi$ in [16], in particular for the case of an infinite-mean $\xi$. In the latter, the results are stated under the assumption of $\xi$ having a probability atom at $1$; however, a straightforward adaptation allows us to present their results without such restriction.
Proposition 2.1 (Theorem 1.1 in [16]).
If $\sum_{j=1}^{n}{\mathbf{P}}\left(\xi\geq j\right)\sim\ell(n)$ for some function $\ell$ slowly varying function at infinity then, as $n\to\infty$, ${\mathbf{E}}\left[M_{n}\right]\sim n/\ell(n)$ and, in probability,
$\frac{M_{n}}{{\mathbf{E}}\left[M_{n}\right]}\stackrel{{\scriptstyle p}}{{\longrightarrow}}1.$
Proposition 2.2 (Theorem 1.4 in [16]).
If ${\mathbf{P}}\left(\xi\geq n\right)\sim\frac{\ell(n)}{n}$ for some function $\ell(n)$ slowly varying at infinity and ${\mathbf{E}}\left[\xi\right]=\infty$, the normalization
$\frac{M_{n}-b(n)}{a(n)}$
converges in distribution to a Cauchy random variable whenever the following conditions hold for positive functions $a(x),b(x)$ and $c(x)$:
  • i)
    $\lim_{x\to\infty}x{\mathbf{P}}\left(\xi\geq c(x)\right)=1$,
  • ii)
    $b(\phi(x))\sim\phi(b(x))\sim x$ where
[…]
  • iii)
    $a(x)\sim x^{-1}b(x)c(b(x))$.
The proofs of Propositions 2.1 and 2.2 rely on a coupling of the process $R^{(n)}_{i}$ with the classic random walk $S_{i}=\sum_{j=1}^{i}\xi_{j}$. Observe that the jumps for $S_{i}$ correspond precisely with the independent variables $\{\xi_{i}:i\in\mathbb{N}\}$; in contrast the jumps of $R^{(n)}_{i}$, being conditioned to not exceed the barrier at $n$, lose both the independence and identical distribution property.
The limiting distribution of $N_{n}:=\inf\{i\in\mathbb{N}:S_{i}\geq n\}$ is well understood and it depends on the tails of $\xi$; moreover, $N_{n}$ may be interpreted as the number of jumps of $S_{i}$ before exceeding the barrier at $n$. The key idea in [15, 16] is to prove that the difference $M_{n}-N_{n}$, under the natural coupling of $(R^{(n)}_{i},S_{i})$, is negligible with respect to the first order of $N_{n}$.
Other related work on the domain of attraction of $\alpha$-stable distributions are the following. Erickson studies the asymptotic behavior of the spent and residual times associated with a renewal process [10]. Geluk and Haan derive the theory of stable probability distributions and their domains of attraction via Fourier transforms [13]. Berger surveys the case of random walks in the Cauchy domain of attraction [3].
2.1 Uniform random cuts
Consider $T_{n}$ a random recursive tree on $n$ vertices and perform a random cut; then the number of vertices removed is given by $\xi^{(n)}$, with
${\mathbf{P}}\left(\xi^{(n)}=k\right)=\frac{n}{(n-1)k(k+1)},$ (3)
for $k\in[n-1]$; see e.g. [18]. On the other hand, let $\xi$ be defined, for $k\in\mathbb{N}$, by
${\mathbf{P}}\left(\xi=k\right)=(k(k+1))^{-1}.$ (4)
Then ${\mathbf{P}}\left(\xi\geq n\right)=n^{-1}$ and the distribution of $\xi$ conditionally given $\xi<n$, coincides with the distribution of $\xi^{(n)}$. It follows that $M_{n}$ corresponds precisely to $X_{n}$, the number of uniform random cuts required to isolate the root of $T_{n}$.
The known properties of $X_{n}$ may be recovered using Propositions 2.1 and 2.2. Namely, as $n$ goes to infinity, ${\mathbf{E}}\left[X_{n}\right]\sim n(\ln n)^{-1}$, and in probability,
$\frac{X_{n}}{{\mathbf{E}}\left[X_{n}\right]}\stackrel{{\scriptstyle p}}{{\longrightarrow}}1;$ (5)
moreover,
$\frac{(\ln n)^{2}}{n}X_{n}-\ln n-\ln(\ln n)$ (6)
converges in distribution to a Cauchy random variable.
3 Analysis of the degree-biased cutting process
Consider a random recursive tree $T_{n}$, $n\geq 2$. In what follows, we let $v_{1}v_{2}$ be a uniformly chosen edge in $T_{n}$ so that, after performing a degree-biased cut to $T_{n}$, we obtain trees $T_{n}^{(v_{1})}$ and $T_{n}^{root}$. The following proposition computes the joint distribution of their relabeled versions.
Lemma 3.1.
Let $T_{n}$ be a random recursive tree of size $n$ and $T_{n}^{root}$, $T_{n}^{(v_{1})}$ be the resulting subtrees of a degree-biased cut. For $2\leq k<n$, $t^{\prime}\in I_{k}$ and $t^{\prime\prime}\in I_{n-k}$ we have that
${\mathbf{P}}\left(\Phi(T_{n}^{(v_{1})})=t^{\prime},\Phi(T_{n}^{root})=t^{\prime\prime}\right)=\dfrac{nd_{t^{\prime}}(1)}{(n-1)(k+1)!(n-k-1)!};$ (7)
while, for $k=n$ and $t^{\prime}\in I_{n}$,
${\mathbf{P}}\left(\Phi(T_{n}^{(v_{1})})=t^{\prime}\right)=\frac{d_{t^{\prime}}(1)}{(n-1)(n-1)!}.$ (8)
Proof.
Note that, conditionally given $v_{1}v_{2}$, we have either $v_{1}=1$, so that $\Phi(T_{n}^{(v_{1})})=T_{n}$ and $T_{n}^{root}=\emptyset$, or $v_{1}>1$ which implies that $T_{n}=T_{n}^{root}\cup T_{n}^{(v_{1})}\cup\{v_{0}v_{1}\}$ and $2\leq|T_{n}^{(v_{1})}|<n$.
For the first case, let $t,t^{\prime}\in I_{n}$, then
${\mathbf{P}}\left(T_{n}^{(v_{1})}=t^{\prime}|T_{n}=t\right)={\mathbf{P}}\left(v_{1}=1|T_{n}=t\right){\mathbf{1}}_{[t^{\prime}=t]}.$
The probability on the right side, given that $t=t^{\prime}$, is precisely the ratio between the degree of the root of $t^{\prime}$ and the number of edges. Therefore
${\mathbf{P}}\left(\Phi(T_{n}^{(v_{1})})=t^{\prime}\right)=\sum_{t\in I_{n}}{\mathbf{P}}\left(T_{n}^{(v_{1})}=t^{\prime}|T_{n}=t\right){\mathbf{P}}\left(T_{n}=t\right)=\frac{d_{t^{\prime}}(1)}{n-1}{\mathbf{P}}\left(T_{n}=t^{\prime}\right).$
Furthermore, $|I_{n}|=(n-1)!$, so that ${\mathbf{P}}\left(T_{n}=t^{\prime}\right)=\frac{1}{(n-1)!}$ and (8) is obtained.
Let $2\leq k<n$, $t^{\prime}\in I_{k}$ and $t^{\prime\prime}\in I_{n-k}$. Similarly, it suffices to verify that
${\mathbf{P}}\left(\Phi(T_{n}^{(v_{1})})=t^{\prime},\Phi(T_{n}^{root})=t^{\prime\prime}|T_{n}=t\right)=\frac{d_{t^{\prime}}(1)}{n-1}{\mathbf{1}}_{[t\in C_{t^{\prime}}]};$ (9)
where $C_{t^{\prime}}$ is the set of increasing trees $t$ of size $n$ such that there is a vertex set $\{v_{0},\dots,v_{k}\}$ such that i) $t^{(v_{1})}$ has vertex set $\{v_{1},\dots,v_{k}\}$, ii) $\Phi(t^{(v_{1})})=t^{\prime}$, iii) $v_{0}v_{1}\in t$ and iv) $\Phi(t\setminus t^{\prime})=t^{\prime\prime}$ (in particular, the vertex set of $t\setminus t^{\prime}$ is $[n]\setminus\{v_{1},\dots,v_{k}\}$).
The above conditions imply that $|C_{t^{\prime}}|=\binom{n}{k+1}$. To see this, let $L=\{v_{0},\dots,v_{k}\}\subset[n]$ and $W=\{w_{1},\ldots,w_{n-k}\}=[n]\setminus\{v_{1},\ldots,v_{k}\}$ satisfy $v_{i}<v_{i+1}$ and $w_{j}\leq w_{j+1}$ for all $i,j$. Denote by $t^{\prime}_{L}$ the tree obtained after relabelling each vertex $i\in t^{\prime}$ with vertex $v_{i}$; similarly, let $t^{\prime\prime}_{W}$ the tree obtained after relabelling each vertex $j\in t^{\prime\prime}$ with vertex $w_{j}$. It is clear that each $t\in C_{t^{\prime}}$ has the form $t=t^{\prime}_{L}\cup t^{\prime\prime}_{W}\cup\{v_{0}v_{1}\}$ for some set $L\subset[n]$ of size $k+1$.
To verify (9) we observe that the probability vanishes whenever $t\notin C_{t^{\prime}}$. On the other side, conditionally given that $t=t^{\prime}_{L}\cup t^{\prime\prime}_{W}\cup\{v_{0}v_{1}\}\in C_{t^{\prime}}$, the event $\{\Phi(T_{n}^{(v_{1})})=t^{\prime},\Phi(T_{n}^{root})=t^{\prime\prime}\}$ is equivalent to choosing one of the $d_{t^{\prime}}(1)$ edges that are children of $v_{1}$ in $t$, this probability is uniform over $t\in C_{t^{\prime}}$. Finally, (9) together with ${\mathbf{P}}\left(T_{n}=t\right)=\frac{1}{(n-1)!}$ yields (7). ∎
The joint distribution of $(T_{n}^{(v_{1})},T_{n}^{root})$ depends only on the size and degree of the tree that is to be cut and not the tree that remains. Thus, conditionally given the size of $|T_{n}^{root}|=\ell$, we have that $T_{n}^{root}$ is equally likely on $I_{\ell}$, so that Proposition 1.2 is established. A similar argument, together with an averaging of $d_{t^{\prime}}(1)$ over $t^{\prime}\in I_{k}$, yields Proposition 1.3.
Proofs of Propositions 1.2 and 1.3.
Let $1<k\leq n$ and $\ell=n-k$, we identify ${\mathbf{E}}\left[d_{T_{k}}(1)\right]$ with the following expression
$\sum_{t^{\prime}\in I_{k}}\frac{d_{t^{\prime}}(1)}{(k-1)!}={\mathbf{E}}\left[d_{T_{k}}(1)\right]=H_{k-1}.$ (10)
We compute the marginal distribution of $\Phi(T_{n}^{root})$. If $\ell\geq 1$, using (7) and (10), we get for any $t^{\prime\prime}\in I_{\ell}$,
${\mathbf{P}}\left(\Phi(T_{n}^{root})=t^{\prime\prime}\right)=\sum_{t^{\prime}\in I_{k}}\frac{nd_{t^{\prime}}(1)}{(n-1)(k+1)!(n-k-1)!}=\frac{nH_{k-1}}{(n-1)(k+1)(k)(n-k-1)!};$
note that the probabilities above depend only on $k=n-|t^{\prime\prime}|$, establishing Proposition 1.2. On the other side, since $|I_{\ell}|=(\ell-1)!=(n-k-1)!$ we have
${\mathbf{P}}\left(|T_{n}^{root}|=\ell\right)=\sum_{t^{\prime\prime}\in I_{\ell}}{\mathbf{P}}\left(\Phi(T_{n}^{root})=t^{\prime\prime}\right)=\frac{nH_{k-1}}{(n-1)(k+1)k};$
this establishes the first term of (1) since $|T_{n}^{root}|=\ell$ if and only if $|T_{n}^{(v_{1})}|=k$. Finally, from (8) and (10), we get the case $k=n$,
${\mathbf{P}}\left(|T_{n}^{(v_{1})}|=n\right)=\sum_{t^{\prime}\in I_{n}}\frac{nd_{t^{\prime}}(1)}{(n-1)(n-1)!}=\frac{nH_{n-1}}{n-1}$
completing the proof of Proposition 1.3. ∎
Proposition 3.2.
Let $K_{0}=K_{1}=0$ and $K_{n}$, $n\geq 2$, be as in Definition 1.4. Then ${\mathbf{P}}\left(K_{2}=1\right)={\mathbf{P}}\left(K_{3}=1\right)=1$ and for $n\geq 4$,
${\mathbf{P}}\left(K_{n}=1\right)$ $=\dfrac{H_{n-2}}{(n-1)^{2}}+\dfrac{H_{n-1}}{n-1},$ (11)
${\mathbf{P}}\left(K_{n}=j\right)$ $=\sum_{k=2}^{n-2}\dfrac{nH_{k-1}}{k(k+1)(n-1)}{\mathbf{P}}\left(K_{n-k}=j-1\right);$ $j\geq 2.$ (12)
Proof.
Let $T_{n}$ be a random recursive tree and let $T^{(v_{1})}_{n}$ be the tree defined after one degree-biased cut of $T_{n}$ and $K_{n}$ be the number of cuts necessary to destroy $T_{n}$.
If the size of $T^{(v_{1})}_{n}$ is either $n-1$ or $n$, then $T_{n}$ is instantly destroyed and $K_{n}=1$. By (1) in Proposition 1.3, we obtain (11) for $n\geq 2$; in particular, $K_{2}=K_{3}=1$ almost surely.
Let $n\geq 4$ and $j\geq 2$. By the splitting property, Proposition 1.2, we have
${\mathbf{P}}\left(K_{n}=j\right)=\sum_{k=2}^{n-2}{\mathbf{P}}\left(|T_{n}^{(v_{1})}|=k\right){\mathbf{P}}\left(K_{n-k}=j-1\right);$
indeed, the cutting procedure after one cut continues on a tree of size $n-k$, conditionally given $|T^{(v_{1})}_{n}|=k$ regardless of the labels in $T_{n}\setminus T^{(v_{1})}_{n}$. Using (1) in the expression above, we obtain (12). ∎
3.1 Harmonic-biased random walk with a barrier
As we mentioned before when $\xi$ has distribution as in (4), the process $(R^{(n)}_{i})_{i\geq 0}$ corresponds precisely to the number of jumps of the classic cutting process of random recursive trees. In the case of the degree-biased cutting, conditionally given a tree of size $n$, the size of the cut $\zeta^{(n)}$ follows the distribution ${\mathbf{P}}\left(\zeta^{(n)}=k\right)=f_{n}(k)n/(n-1)$ where
$f_{n}(k):=\frac{H_{k-1}}{k(k+1)}{\mathbf{1}}_{[2\leq k<n]}+\frac{H_{k-1}}{k}{\mathbf{1}}_{[k=n]}.$ (13)
Unfortunately, there is no distribution $\zeta$ of positive integers such that, for $2\leq k\leq n$,
${\mathbf{P}}\left(\zeta=k|\zeta\leq n\right)={\mathbf{P}}\left(\zeta^{(n)}=k\right)$ (14)
To see this, write $p_{k}:={\mathbf{P}}\left(\zeta=k\right)$. Rearranging the terms of (14), we need
$\frac{n}{n-1}\sum_{i=2}^{n}p_{i}=\frac{p_{k}}{f_{n}(k)}$ (15)
to hold for all $2\leq k\leq n$. Since the left side does not depend on $k$ and converges to one, we infer that it is necessary that $p_{k}=\frac{H_{k-1}}{k(k+1)}$ for all $k\geq 2$. However, in such case, (15) fails to hold when $k=n$.
Remark 3.3.
The probability that the degree-biased cutting procedure is destroyed in one step has a relatively large mass coming from an instant destruction of the tree (case $k=n$ in (13)). Thus, $\zeta^{(n)}$ has anomalous behaviour at its largest atom $n$, preventing an accurate approximation by an unbounded random variable $\zeta$.
In what follows we let $(R^{(n)}_{i})_{i\geq 0}$ be the random walk with a barrier on $n$ with jump distribution $\zeta$ given by ${\mathbf{P}}\left(\zeta=k\right)=\frac{H_{k-1}}{k(k+1)}$ and let $J_{n}$ be the number of jumps of the process $(R^{(n+1)}_{i})_{i\geq 0}$; see Section 1.2. Then $\zeta$ is our best jump distribution approximation for the number of vertices discarded after each degree-biased cut. The role of $\zeta$ is analogous to $\xi$, defined (3), for the classic random cutting process of $T_{n}$.
Let $\tau=\inf\{i\in\mathbb{N}_{0}:R^{(n+1)}_{i}>0\}$. By the strong Markov property, the process $(R^{(n+1)}_{\tau+j})_{j\geq 0}$ is independent of $(R^{(n+1)}_{i})_{0\leq i\leq\tau}$ given $R^{(n+1)}_{\tau}$. In particular, conditionally given $R^{(n+1)}_{\tau}=k$, $1<k\leq n$, $(R^{(n+1)}_{\tau+j})_{j\geq 0}$ has the same distribution as $(R^{(n-k+1)}_{j})_{j\geq 0}$. Let $J_{0}=J_{1}=0$. Similarly to Proposition 3.2, for $j\geq 2$,
${\mathbf{P}}\left(J_{n}=j\right)=\sum_{k=2}^{n-2}{\mathbf{P}}\left(\zeta=k\,|\,\zeta\leq n\right){\mathbf{P}}\left(J_{n-k}=j-1\right).$ (16)
We now prepare for the proofs of Theorems 1.5 and 1.6. Using a telescopic argument, we have for $i\geq 0$,
$\sum_{k=i+1}^{n}\frac{1}{k(k+1)}=\sum_{k=i+1}^{n}\frac{1}{k}-\frac{1}{k+1}=\frac{1}{i+1}-\frac{1}{n+1}.$ (17)
In turn, for $n\geq 2$,
${\mathbf{P}}\left(\zeta\leq n\right)=\sum_{k=2}^{n}\sum_{i=1}^{k-1}\dfrac{1}{ik(k+1)}=\sum_{i=1}^{n-1}\dfrac{1}{i}\sum_{k=i+1}^{n}\dfrac{1}{k(k+1)}=1-\frac{1}{n}-\frac{H_{n-1}}{n+1}.$ (18)
This verifies, in passing, that $\zeta$ is a random variable on $\mathbb{N}$. In addition, note that
$\left(1-\frac{1}{n}-\frac{H_{n-1}}{n+1}\right)^{-1}=\frac{n(n+1)}{n^{2}-nH_{n-1}-1}\geq\frac{n+1}{n-H_{n-1}}\geq\frac{n}{n-1}.$
Consequently, for $2\leq k<n$,
${\mathbf{P}}\left(\zeta=k|\zeta\leq n\right)=\frac{H_{k-1}}{k(k+1)}\left(1-\frac{1}{n}-\frac{H_{n-1}}{n+1}\right)^{-1}\geq{\mathbf{P}}\left(\zeta^{(n)}=k\right).$ (19)
Proof of Theorem 1.6.
For $n\leq 3$, $K_{n}=J_{n}$ almost surely. For $n\geq 4$, it suffices to verify that ${\mathbf{P}}\left(J_{n}\geq j\right)\geq{\mathbf{P}}\left(K_{n}\geq j\right)$ for all $j\in\mathbb{N}$. Clearly ${\mathbf{P}}\left(J_{n}\geq 1\right)={\mathbf{P}}\left(K_{n}\geq 1\right)=1$, so henceforth we assume that $j\geq 2$. From the expression in (16), by changing the order of the sums we get
${\mathbf{P}}\left(J_{n}\geq j\right)$ $=\sum_{i=j}^{\infty}\sum_{k=2}^{n-2}{\mathbf{P}}\left(\zeta=k\,|\,\zeta\leq n\right){\mathbf{P}}\left(J_{n-k}=i-1\right)$
$=\sum_{k=2}^{n-2}{\mathbf{P}}\left(\zeta=k\,|\,\zeta\leq n\right){\mathbf{P}}\left(J_{n-k}\geq j-1\right);$ (20)
analogously, using (12),
${\mathbf{P}}\left(K_{n}\geq j\right)$ $=\sum_{k=2}^{n-2}{\mathbf{P}}\left(\zeta^{(n)}=k\right){\mathbf{P}}\left(K_{n-k}\geq j-1\right).$ (21)
The proof of Theorem 1.6 is complete by induction on $j\geq 2$ and $n\geq 4$ using (19)–(21). ∎
Proof of Theorem 1.5.
From (18) we have that the tails of $\zeta$ satisfy
${\mathbf{P}}\left(\zeta\geq n\right)=1-{\mathbf{P}}\left(\zeta\leq n-1\right)=\frac{1}{n-1}+\frac{H_{n-2}}{n}=\frac{\ln n}{n}(1+o(1)).$ (22)
The conditions of Proposition 2.1 are satisfied for
$g(n)=\sum_{j=1}^{n}{\mathbf{P}}\left(\zeta\geq j\right)\sim\frac{(\ln n)^{2}}{2}.$
Indeed, $g(n)$ is slowly varying at infinity; see Lemma A.1. Therefore, ${\mathbf{E}}\left[J_{n}\right]\sim 2n(\ln n)^{-2}$ and $\frac{J_{n}}{{\mathbf{E}}\left[J_{n}\right]}$ converges in probability to one.
We now verify that the conditions of Proposition 2.2 are satisfied for $a(x)=4x(\ln x)^{-3}$, $b(x)=2x(\ln x)^{-2}$ and $c(x)=x\ln x$.
First, substitution of $n=x\ln x$ in (22) yields
$\lim_{x\to\infty}x{\mathbf{P}}\left(\zeta\geq x\ln x\right)=\lim_{x\to\infty}\frac{x(\ln x+\ln\ln x)}{x\ln x}(1+o(1))=1.$
Again, by using (22) in the definition of $\phi(x)$ and
[…]
we have $\phi(x)\sim\frac{x(\ln x)^{2}}{2}$ since
[…]
where we used that $\ln(x\ln x)\sim\ln x$. Similarly, the fact that $\ln x$ is a slowly varying function implies that $\ln 2x-2\ln\ln x\sim\ln x+(2\ln\ln x-\ln 2)\sim\ln x$. Therefore
$\phi(b(x))\sim\frac{x(\ln 2x-2\ln\ln x)^{2}}{(\ln{x})^{2}}\sim x\sim\frac{x(\ln x)^{2}}{(\ln x+(2\ln\ln x-\ln 2))^{2}}\sim b(\phi(x));$
whereas,
$x^{-1}b(x)c(b(x))=\dfrac{4x(\ln{x}-2\ln{\ln{x}}-\ln 2)}{(\ln{x})^{4}}\sim a(x).$
Proposition 2.1 then implies that
$\frac{J_{n}-b(n)}{a(n)}=\frac{J_{n}}{4n}(\ln n)^{3}-\frac{\ln n}{2}$
converges in distribution to a Cauchy random variable. Finally, [12, Lemma 1, VIII.2, p. 253] implies that such convergence also holds for the normalization given by
$a^{\prime}_{n}=\frac{n}{(\ln{n})^{3}}\text{ and }b^{\prime}_{n}=\frac{n}{(\ln{n})^{2}},$
which recovers the statement in Theorem 1.5. ∎
4 Related models: cut-tree and coalescent process
The cut-tree is a binary tree that encodes the structure of all connected components of $T_{n}$ as the edges are removed from the graph. Vertices in the cut-tree are labelled by the vertex set of the connected components which arise during the (percolation) process; in particular the leaves of the cut-tree are singletons representing the vertices of $T_{n}$. More precisely, the root is labelled $[n]$ and its two children are labelled $V_{1}$ and $V_{2}$ and correspond to the vertex sets of the subtrees obtained after the first edge-removal of the process; we then proceed in a similar fashion on both subtrees.
The cut-tree was initially defined in the context of Galton-Watson trees; for example, the cut-tree of a critical Galton–Watson tree with finite variance and conditioned to have size $n$ converges as $n\to\infty$ to a Brownian continuum random tree (CRT) in the weak sense induced by the Gromov-Prokhorov topology [5]. For the case of random recursive trees, Bertoin showed that, rescaled by the factor $n^{-1}\ln n$, $\mathrm{Cut}(T_{n})$ converges in the Gromov-Hausdorff-Prokhorov topology to the line of unit length, [4]. Previously known properties of the edge-removal process were recovered directly from the cut-tree. For example, the time to isolate a given vertex $v$ (not necessarily the root) corresponds to the height of $\{v\}$ in $Cut(T_{n})$.
There is also a natural coalescent process associated with the cutting procedure of a random tree. A coalescent process is a Markov process that takes its values in the set of partitions of $\mathbb{N}$ and it is generally described by its restrictions to the set $[n]$, for $n\in\mathbb{N}$.
The general procedure for obtaining a $n$-coalescent from a cutting procedure is the following. At any given step we have a tree whose vertex set $\{v_{1},v_{2},\ldots,v_{m}\}$ are the blocks of a partition of $[n]$; moreover, the blocks are listed in increasing order of least elements and all paths starting from the root $v_{1}$ are increasing. Select a random vertex $v$ (which in the cutting process would be removed) and add the labels of the removed vertices into the label $v$; in other words, coalesce all the blocks of the subtree of $v$, inclusive, to form a tree on a new label-set. The $n$-coalescent is the process of the partition of $[n]$ corresponding to the label set; starting from a tree with all singletons.
Goldschmidt and Martin first proposed the connection between cutting procedures and coalescent processes [14]. They identified that the coalescent starting from a random recursive tree $T_{n}$ defines the Bolthausen Sznitman coalescent on $\mathbb{N}$. Conditionally given that there are $b$ blocks in the partition, for $2\leq k\leq b\leq n$, the rate at which any given set of $k$ labels coalesce is
$\frac{(b-k)!(k-2)!}{(n-1)!}.$ (24)
Procedures to obtain beta-coalescent use uniformly random binary trees and Galton-Watson trees with offspring distribution in the domain of attraction of a stable law of index $\alpha\in[1/2,1)$ [1].
A similar process is studied in [20] for plane-oriented recursive trees (also identified as preferential attachment trees): Select a vertex $v_{1}$ uniformly at random. If $v_{1}$ is a leaf, do nothing. Otherwise, select one of its successors, say $v_{2}$, uniformly at random. Proceed as the process described above with the edge $v_{1}v_{2}$. In this case, the resulting coalescent process is related to an arcsine coalescent; however, the finite coalescent processes are not consistent, that is, these cannot be extended to a coalescent process on $\mathbb{N}$.
4.1 Degree-biased versions and open problems
We define the degree-biased coalescent process as follows: start with $T_{n}$ a random recursive tree on $[n]$ and associate an independent exponential random variable with mean $1$ to each edge. This random variable is the time when the edge is selected to perform the first step of the degree-biased cutting process.
If the edge $\ell_{i}\ell_{j}$ is selected and $i>1$, then there exists $\ell_{h}$ such that $\ell_{h}\ell_{i}$ is an edge on $T_{n}$. The labels in the subtree $T^{(\ell_{i})}$ are instantly added to the label of the vertex $\ell_{h}$. That is, the new set of labels replaces $\ell_{h}$ with $\ell_{h}\cup\{s:s\in T_{n}^{(\ell_{i})}\}$.
The proof of the following proposition is based on Proposition 1.3 and the ideas in Lemma 3.1.
Proposition 4.1.
The first coalescing event corresponding to the degree-biased cutting of a random recursive tree of size $n$ merges any given set of $k$ labels at rate
$\lambda^{\prime}_{n,k}$ $=\dfrac{(n-k)!(k-2)!}{(n-1)!}H_{k-2}\quad\text{for $3\leq k<n$},$ (25)
$\lambda^{\prime}_{n,n}$ $=H_{n-1}+\dfrac{H_{n-2}}{n-1}.$ (26)
Proof.
We start with the case $k=n$ corresponding to having all singletons merged in one step; from the degree-biased cutting procedure, this is equivalent to having the process end in exactly one step. Namely, either the tree is instantly destroyed (an edge adjacent to the root is selected in the cutting process) or the root has exactly one vertex attached, vertex 2, and the selected edge $v_{1}v_{2}$ has $v_{1}=2$. In other words, the probability that all blocks coalescent corresponds to the probability that $|T_{n}^{(v_{1})}|\in\{n-1,n\}$. Since the next coalescent event is the minimum of $n-1$ exponential random variables with rate 1, the total rate of instant coagulation is, by Proposition 1.3,
$\lambda^{\prime}_{n,n}=(n-1){\mathbf{P}}\left(|T_{n}^{(v_{1})}|\in\{n-1,n\}\right)=H_{n-1}+\frac{H_{n-2}}{n-1}.$
For the case $3\leq k<n$, let $L=\{v_{0},v_{1},\ldots,v_{k-1}\}$ be a set of $k$ vertices in $T_{n}$ such that $v_{0}<v_{1}<\cdots<v_{k-1}$. Then, the rate $\lambda_{L}$ at which elements in $L$ coalesce is given by
$\lambda_{L}=(n-1){\mathbf{P}}\left(V(T_{n}^{(v_{1})})=L\right).$
For each $t^{\prime}\in I_{k-1}$, let $C_{L,t^{\prime}}$ be the set of increasing trees $t$ of size $n$ such that i) $t^{(v_{1})}$ has vertex set $\{v_{1},\dots,v_{k-1}\}$, ii) $\Phi(t^{(v_{1})})=t^{\prime}$ and iii) $v_{0}v_{1}\in t$. By a similar argument as that in the proof of Lemma 3.1, we have
${\mathbf{P}}\left(V(T_{n}^{(v_{1})})=L\right)$ $=\sum_{t^{\prime}\in I_{k-1}}\sum_{t\in C_{L,t^{\prime}}}{\mathbf{P}}\left(T_{n}=t,\Phi(T_{n}^{(v_{1})})=t^{\prime}\right)$ $=\sum_{t^{\prime}\in I_{k-1}}\sum_{t\in C_{L,t^{\prime}}}\frac{d_{t^{\prime}}(1)}{(n-1)(n-1)!};$
using $|C_{L,t^{\prime}}|=|I_{n-k+1}|=(n-k)!$ together with (10), we get
$\lambda_{L}=\frac{(n-k)!}{(n-1)!}\sum_{t^{\prime}\in I_{k-1}}d_{t^{\prime}}(1)=\frac{(n-k)!(k-2)!}{(n-1)!}H_{k-2}.$
The rate $\lambda_{L}$ depends only on the size of $L$; completing the proof of (25). ∎
As we mentioned before, these coalescents can not be extended to a $\Lambda$-coalescent process on $\mathbb{N}$ since their rates do not satisfy the consistency equations $\lambda^{\prime}_{n,k}=\lambda^{\prime}_{n+1,k}+\lambda^{\prime}_{n+1,k+1}$. For example, the following expression does not vanish,
$\lambda^{\prime}_{n,n}+\lambda^{\prime}_{n,n-1}-\lambda^{\prime}_{n-1,n-1}$ $=H_{n-1}-H_{n-2}+\frac{H_{n-2}}{n-1}+\frac{H_{n-3}}{(n-1)(n-2)}-\frac{H_{n-3}}{n-2};$
after a careful rearrangement and cancellation of terms, it simplifies to
$\frac{1}{n-1}+\frac{H_{n-2}}{n-1}-\frac{H_{n-3}}{n-1}=\frac{1}{n-1}+\frac{1}{(n-1)(n-2)}=\frac{1}{n-2}>0.$
Appendix A A slowly varying function at infinity
Recall $g(n)$ defined in the proof of Theorem 1.5 in Section 3.1. Together with (22),
$g(n)=\sum_{j=1}^{n}{\mathbf{P}}\left(\zeta\geq j\right)=1+\sum_{j=2}^{n}\left(\frac{1}{j-1}+\frac{H_{j-2}}{j}\right).$
Lemma A.1.
The function $g(n)$ is slowly varying at infinity and $g(n)\sim\frac{(\ln n)^{2}}{2}$.
Proof.
Rearranging the terms of $g(n)$ we get
$g(n)$ $=1+H_{n-1}+\sum_{j=3}^{n}\frac{H_{j}}{j}-\sum_{j=3}^{n}\left(\frac{1}{j(j-1)}+\frac{1}{j^{2}}\right)=H_{n-1}+c+\varepsilon_{n}+\sum_{j=1}^{n}\frac{H_{j}}{j};$
where $c=\frac{3-\pi^{2}}{6}$ and $\varepsilon_{n}\to 0$ as $n\to\infty$. Since the class of slowly varying functions is closed under addition, it remains to prove that $\sum_{j=1}^{n}\frac{H_{j}}{j}$ is slowly varying. To see this note that
$\sum_{j=1}^{n}\frac{H_{j}}{j}$ $=\sum_{j=1}^{n}\sum_{i=1}^{j}\frac{1}{ij}=\sum_{i=1}^{n}\frac{1}{i}\sum_{j=i}^{n}\frac{1}{j}=\sum_{i=1}^{n}\frac{1}{i}(H_{n}-H_{i-1});$
rewriting $H_{i-1}=H_{i}-\frac{1}{i}$ and regrouping terms yields
$\sum_{j=1}^{n}\frac{H_{j}}{j}=\frac{1}{2}\left(H_{n}^{2}+\sum_{j=1}^{n}\frac{1}{j^{2}}\right).$
This completes the proof as both terms on the right side are slowly varying and
$g(n)=\left(\sum_{j=1}^{n}\frac{H_{j}}{j}\right)(1+o(1))\sim\frac{(\ln n)^{2}}{2}.$